# Tiling spaces are covering spaces over irrational tori

### v1

@@ Abstract
We study tiling spaces in the diffeological context. We prove some basic diffeological properties for tiling spaces and analyze two different fiber bundle structures of one-dimensional tiling spaces over the irrational torus. We use the diffeological classification of irrationatl tori which captures their arithmetical escence in order to inherit the diffeological equivalence in the context of tiling spaces.
@@ 1 Introduction
Motivation comes from the naive question: what is the aperiodic analog of the orbifolds that characterize the symmetries of periodic tilings of the plane? Conway’s Magic Theorem [9] classifies the symmetries of periodic tilings of the plane through the orbifolds obtained as quotients $\mathcal{O}=\mathbb{R}^{2}/G$, where $G$ is the symmetry group of the tiling. In the crystallographic case, $G$ has a subgroup of translations isomorphic to $\mathbb{Z}^{2}$. Thus the aforementioned orbifolds are quotients of the torus $\mathbb{T}^{2}$ by a discrete subgroup of $G$. In the aperiodic one-dimensional case, however, one will find that a plausible analog of the translations subgroups $\mathbb{Z}$ could be (a subset of) $\mathbb{Z}+\alpha\mathbb{Z}$, with $\alpha$ an irrational number. So one should look at $\mathbb{T}_{\alpha}:=\mathbb{R}/(\mathbb{Z}+\alpha\mathbb{Z})$, which is quite a strange object, its natural topology is coarse. It is nevertheless non-trivial when endowed with a structure of diffeological space, it is called the irrational torus of slope $\alpha$ and it was extensively studied by Iglesias-Zemmour in [17, 18]. On the other hand, another possible analog for $\mathcal{O}$ could be the tiling space $\Omega$, also known as the continuous hull of a tiling. The tiling space of a periodic tiling is a torus obtained by identification by translations, if further symmetries are to be considered then one will end up with Conway’s orbifolds.
Our approach is to define a structure of diffeological space for $\Omega$ and analyze its relationships with the irrational torus $\mathbb{T}_{\alpha}$. We verify some basic diffeological properties for $\Omega$ and explore two different fiber bundles constructions of $\Omega$ over $\mathbb{T}_{\alpha}$. The first one is a variant of the fiber bundle over the torus with Cantor set fibers first constructed by Williams [36] and later on generalized by Williams himself and Sadun [34]. This variant turns out to be more natural in the sense that there is no need to deform the tiles in order to construct it, so it holds up to stronger equivalence – we include a discussion of different sorts of equivalence in geometrical and combinatorial contexts. This result is stated as
@@
Let $\Omega$ be an unidimensional tiling space of primitive substitution tilings with a rule $\omega$ and expansion factor $\alpha$. Then $\Omega$ is a covering space over the irrational torus $\mathbb{T}_{\alpha}$.
@@
Let $\Omega_{\alpha}$ a tiling space of canonical projection of dimension 2 to 1 and irrational slope $\alpha$. Then $\Omega_{\alpha}$ is a $\mathbb{R}$-principal bundle over the double pointed irrational torus $\mathbb{T}_{\alpha}^{\bullet}$.
@@
$\Omega_{\alpha}$ and $\Omega_{\beta}$ are strong orbit equivalent.
Their associated sturmian spaces $X_{\alpha},X_{\beta}$ are strong orbit equivalent.
$\mathbb{T}_{\alpha}$ and $\mathbb{T}_{\beta}$ are diffeomorphic.
@@
These results imply some other nice properties. On the one hand, the existence of such a bundle allows us to define and compute a fundamental group for tiling spaces thanks to the long exact sequence of homotopy groups – this appears as Corollary 2. On the other, by using previous work on the existence of substitution rules for projection method tilings [27, 16] and the arithmetic implications of the mentioned equivalence, we prove the following
Theorem D.
This kind of problems have been extensively studied from different approaches as topological [24, 6, 22, 29, 33], dynamical [31, 1, 4, 3, 11], and algebraic [8, 10, 12, 13, 35, 25, 30], among many others. Our results are in principle similar in many cases. It is worth mentioning that there are some recent advances about the links of the diffeological theory and the Non-commutative geometry in more general settings [19, 20, 23]. Our work could provide a series of examples for these recent advances.
The rest of the article is organized as follows. In Section 2 we introduce the basic notions of both tiling spaces and diffeology, we provide in each case several definitions and state the results that we use throughout the text. In Section 3 we define and study a diffeological version of tilings spaces, we construct the first of the mentioned bundles $\hat{p}:\Omega\to\mathbb{T}_{\alpha}$ and investigate its structure groupoid, in order to prove Theorem A. Section 4 is devoted to investigate the strong orbital equivalence between tiling spaces and to the proof of the remaining results.
2 Preliminars
@@ 2.1.1 Tilings
    Each $t_{i}$ is a translated copy of some element of a finite set of proto-tiles $\mathcal{P}=\{\tau_{1},\ldots,\tau_{2}\}$.
@@
A tiling $T$ is periodic if its symmetry group has an infinite cyclic subgroup and aperiodic in other case. There are at least three methods for constructing aperiodic tilings: substitution, projection, and local rules. We will consider only the first two classes.
A substitution rule is a function from the set of proto-tiles onto the set of patterns. It consists of a subdivision rule for the proto-tile followed by a homothety, resulting in a pattern of the original set of proto-tiles. By iterating a substitution rule over a finite initial configuration of tiles, arbitrarily large portions of an euclidean space can be covered and thus a tiling can be constructed.
@@ 2.1.2 Tiling spaces
By definition, $\Omega_{T}$ contains every tiling $T^{\prime}$ which is locally the same as $T$, meaning that every patch of $T^{\prime}$ may be found on $T$. By patch we mean a finite and connected union of tiles. In general, a tiling space $\Omega$ is a translation invariant set of tilings which is complete in the metric $d$.
@@
A substitution rule $\omega$ is called primitive if every proto-tile appears in some iteration of the substitution rule applied to any other proto-tile. Primitive substitutions generate an unique LI class, thus every tiling generate the same hull and so we may speak about a tiling space $\Omega_{\omega}$ associated to $\omega$.
A tiling $T$ has finite local complexity (FLC) if for every real number $R>0$ there is a finite number of patches, up to translation, that fit inside the ball of radius $R$; and $T$ is called repetitive if for every pattern $P$ of $T$ there is a radius $R>0$ such that for every $x\in\mathbb{R}^{d}$ the ball $B_{R}(x)$ contains a translated copy of $P$.
@@
Let $T$ be an $FLC$ tiling. Then $\Omega_{T}$ is compact and the following conditions are equivalent:
@@
This fairly large class of tilings produces nice structures on the tiling spaces. The hulls $\Omega$ of such tilings are locally homeomorphic to a cartesian product of an euclidean ball $D$ times a totally disconnected set $K$ (a Cantor set under additional assumptions); they are a rich source of examples for hyperbolic dynamics, they are known to be laminated spaces [15], inverse limits of branched manifolds [2, 32, 7] and fiber bundles over tori [36, 34]. For a deeper introduction to the topology of tiling spaces see [33]. In the following paragraphs, we will state the theorems and briefly explain the intuition behind the last couple of results enlisted.
@@
Intuition behind Theorem 2. The iteration of a substitution rule assures that arbitrarily large euclidean balls can be covered. This procedure can be encoded in the structure of an inverse limit. Regard each iteration of a substitution rule both as a finite pattern and a cylinder set, meaning that the finite pattern also represents all the tilings showing that pattern in some particular position (the canonical choice being around the origin). This cylinder set is identified with a branched manifold in the following way: consider the disjoint union of the set of prototiles and identify two of them by their borders if they appear next to each other somewhere in the tiling. Let $\Omega_{0}$ be the space obtained after the identification. It is a $CW$-complex, and it is actually a branched manifold. Then let $\Omega_{n}$ be the branched manifold obtained by applying to $\Omega_{0}$ the substitution rule $n$ times (see Fig. 1). The theorem by Anderson and Putnam [2] asserts that the tiling space $\Omega$ is topologically conjugate to the inverse limit of ${\Omega_{n},f_{n}}$ where the $f_{n}$ maps consist in forgetting the extra information from $\Omega_{n}$ to fit in $\Omega_{n-1}$. In other words, we can regard $\Omega$ as a subset of the infinite product of branched manifolds $\Pi_{n\in\mathbb{N}}\Omega_{n}$, and
$\Omega=\lim_{\leftarrow i}(\Omega_{n},f_{n})=\{(\bar{x}_{n})\in\Pi_{n\in\mathbb{N}}\Omega_{n}:f_{n+1}(\bar{x}_{n+1})=\bar{x}_{n},\forall n\in\mathbb{N})\}.$
@@
Figure 1: The branched 1-dimensional manifolds $\Omega_{0}$, $\Omega_{1}$ and $\Omega_{2}$ associated to the subtitution rule $a\mapsto b$, $b\mapsto ab$.
@@
The idea for constructing such a bundle uses the previous result and the canonical projection $p_{0}$ from the inverse limit (which can be identified with) $\Omega$ onto the branched manifold $\Omega_{0}$, and then constructing a map from $\Omega_{0}$ which identifies each proto-tile with a torus of the corresponding dimension. The construction of the later map relies on the fact that tiles can be deformed in order for their vertices to have rational coordinates. Such a deformation works only up to homeomorphism of tiling spaces.
Regarding the fiber, in the cases of substitution or projection tilings, it is a Cantor set. Whereas for periodic tilings it is a finite collection of points.
One of the results of this article is the construction of a variation of this bundle in the diffeological setting.
@@ 2.1.3 Equivalence of tiling spaces
  • Homeomorphism, in the usual topological sense. This is the weaker equivalence of this list. Homeomorphisms map translational orbits into translational orbits, but not much more than that.
@@
Topological conjugacy is a rather rigid property among tilings. To exemplify this we reproduce the following example from [33]. Consider a one-dimensional tiling space $\Omega$ coming from the Fibonacci substitution $a\mapsto b$, $b\mapsto ab$, with tile lengths $|a|=1$ and $|b|=\varphi=\frac{1+\sqrt{5}}{2}$. Let $a_{n},b_{n}$ denote the patches (words) obtained by applying the substitution $n$ times to $a$ and $b$, respectively. Let $\Omega^{\prime}$ be a suspension of the Fibonacci subshift with tile lengths $|a^{\prime}|=|b^{\prime}|=(2+\varphi)/(1+\varphi)$, and let $a^{\prime}_{n}$ and $b_{n}^{\prime}$ the analogs of $a_{n},b_{n}$. The spaces $\Omega$ and $\Omega^{\prime}$ are known to be topologically conjugate by a mapping which preserves the sequences of tiles. This happens because the differences in lengths $|a_{n}|-|a^{\prime}_{n}|$ and $|b_{n}|-|b_{n}^{\prime}|$ go to zero as $n$ tends to infinity.
The first kind of equivalence between tiling spaces that we will encounter in sintony with diffeology equivalences is the following:
@@
Two tiling spaces $\Omega$ and $\Omega^{\prime}$ are orbit equivalent if there exist a homeomorphism $h:\Omega\to\Omega^{\prime}$ mapping the orbit $O(T)$ in $\Omega$ onto the orbit $O(h(T))$ in $\Omega^{\prime}$ for every tiling $T\in\Omega$.
@@ 2.1.4 Geometrical and combinatorial aspects of tiling spaces
A cut-and-project set (or CPS) is a triple $(E,H,\mathcal{L})$ where $E\cong\mathbb{R}^{d}$ is the physical space of the tiling, $H$ is a locally compact abelian group called internal space, a lattices $\mathcal{L}\subset\mathbb{R}^{d}\times H$, and the canonical projections $\pi_{1}:\mathbb{R}^{d}\times H\to\mathbb{R}^{d}$ and $\pi_{2}:\mathbb{R}^{d}\times H\to H$ such that $\pi_{1}|_{\mathcal{L}}$ is injective and $\pi_{2}(\mathcal{L})$ is dense in $H$.
@@
$\Lambda(W)=\pi_{1}(\{y\in\mathcal{L}:\pi_{2}(y)\in W\}).$
@@
The tiling space of a $(\mathbb{R}^{d},\mathbb{R}^{n},\mathcal{L})$-CPS is closely related to a torus of dimension $n+d$. We explain further the case of projection from dimension 2 to 1. Consider the line $L_{\alpha}=\{y=\alpha x\}$ with slope the irrational number $\alpha$. Thus, the $(L_{\alpha},\mathbb{R},\mathbb{Z}^{2})$-CPS of canonical projection is an aperiodic tiling $T(\alpha)$ with two tiles labelled with $0,1$ of lengths $\frac{1}{\sqrt{1+\alpha^{2}}},\frac{\alpha}{\sqrt{1+\alpha^{2}}}$ respectively. Observe that each point $p$ in $L_{\alpha}$ corresponds to an element of $O(T(\alpha))$ by the relation
@@
Note that there exist an ambiguity whenever the line $L_{\alpha}$ intersects a point with integer coordinates, say $(0,0)$ for simplicity, as both projections $\pi_{2}(0,1)$ and $\pi_{2}(1,0)$ fall in the border of the canonical window $W$. Selecting either of them give us a different tiling, and these two tilings only differ in the tiles located on their origins, so they are at maximun distance in the tiling metric.
@@
with a double solenoid $\mathcal{S}_{\alpha}^{\prime}=\pi(L_{\alpha})^{(0,1)}\cup\pi(L_{\alpha})^{(1,0)}$ and every translational orbit is represented by the projection $\pi(y=\alpha x+\rho)$ with $\rho\in(0,1)$ in $\mathbb{T}^{2}$. The double solenoid $\mathcal{S}_{\alpha}^{\prime}$ comes from the ambiguity mentioned above. See Section 1.4 in [33], and [12] for a more complete account.
@@
Given a tiling $T=\{t_{i}\}_{i\in\mathbb{Z}}$ of $\mathbb{R}$ we can identify its set of proto-tiles $\{\tau_{j}\}$ with numbers $\{a_{j}\}$ so as to obtain a sequence $S=\{s_{i}\}_{i\in\mathbb{Z}}$ which take values in $\{a_{j}\}$. The sequence $S$ is the combinatorial aspect of the tiling $T$. This kind of sequences, endowed with the shift map
@@
The sturmian word of slope $\alpha$ and intercept $\rho$ can also be encoded as the cutting sequence111 The cutting sequence of a line $L_{\alpha}$ is constructed by writing 1 anytime the line $L_{\alpha}$ crosses a horizontal line between the parallels lines $x=n-1$ and $x=n$ and 0 when it doesn’t cross any horizontal line. of the line $y=\alpha x+\rho$. In terms of their complexity, sturmian sequences are characterized as the words containing exactly $n+1$ different subwords of lenght $n$. The literature for sturmian sequences is vast (see for example [26]). We will only focus on the types of equivalences among sturmian sequences that turn out to be relevant in our context.
@@
Two sturmian spaces $(X,\sigma)$ and $(Y,\sigma)$ are orbit equivalent if there exist a homeomorphism $h:X\to Y$ mapping the orbit $O_{\sigma}(x)$ of $x$ under $\sigma$ in $X$ onto the orbit $O_{\sigma}(h(x))$ under $\sigma$ of $h(x)$ in $Y$ for every $x\in X$.
@@
Two sturmian spaces $(X,\sigma)$ and $(Y,\sigma)$ are strong orbit equivalent if they are orbit equivalent and the cocycle map has an unique point of discontinuity.
@@
    $X_{\alpha}$ and $X_{\beta}$ are strong orbit equivalent;

### v3

@@ Abstract
We study tiling spaces in the diffeological context. We prove some basic diffeological properties for tiling spaces and analyze two different fiber bundle structures of one-dimensional tiling spaces over the irrational torus. We use the diffeological classification of irrational tori which captures their arithmetical escence in order to inherit the diffeological equivalence in the context of tiling spaces.
@@ 1 Introduction
Our motivation comes from the naive question: what is the aperiodic analog of the orbifolds that characterize the symmetries of periodic tilings of the plane? Conway’s Magic Theorem [9] classifies the symmetries of periodic tilings of the plane through the orbifolds obtained as quotients $\mathcal{O}=\mathbb{R}^{2}/G$, where $G$ is the symmetry group of the tiling. In the crystallographic case, $G$ has a subgroup of translations isomorphic to $\mathbb{Z}^{2}$. Thus the aforementioned orbifolds are quotients of the torus $\mathbb{T}^{2}$ by a discrete subgroup of $G$. In the aperiodic one-dimensional case, however, one will find that a plausible analog of the translations subgroups $\mathbb{Z}$ could be (a subset of) $\mathbb{Z}+\alpha\mathbb{Z}$, with $\alpha$ an irrational number. So one should look at $\mathbb{T}_{\alpha}:=\mathbb{R}/(\mathbb{Z}+\alpha\mathbb{Z})$, which is quite a strange object, its natural topology is coarse. It is nevertheless non-trivial when endowed with a structure of diffeological space, it is called the irrational torus of slope $\alpha$. It was first studied by Iglesias-Zemmour and Donato in [20], see also [21, 18]. On the other hand, another possible analog for $\mathcal{O}$ could be the tiling space $\Omega$, also known as the continuous hull of a tiling. For instance, the tiling space of a periodic tiling is a torus obtained by identifying translations. If further symmetries are to be considered then one will end up with Conway’s orbifolds.
Our approach is to define a diffeological space structure on $\Omega$ and analyze its relationships with the irrational torus $\mathbb{T}_{\alpha}$. We verify some basic diffeological properties for $\Omega$ and explore two different fiber bundle constructions of $\Omega$ over $\mathbb{T}_{\alpha}$. The first one is a variant of the fiber bundle over the torus with Cantor set fibers first constructed by Williams [38] and later on generalized by Williams himself and Sadun [36]. This variant turns out to be more natural in the sense that there is no need to deform the tiles in order to construct it, so it holds up to stronger equivalence – we include a discussion of different sorts of equivalence in geometrical and combinatorial contexts. This result is stated as
@@
Let $\Omega$ be a space of repetitive tilings of $\mathbb{R}^{d}$. Then $\Omega$ is a covering space over $\mathbb{R}^{d}/R_{v}=\mathbb{T}_{v}$, where $R_{v}$ is the return module of a vertex $v$.
@@
Let $\Omega_{\alpha}$ be a tiling space of canonical projection of dimension 2 to 1 and irrational slope $\alpha$. Then $\Omega_{\alpha}$ is an $\mathbb{R}$-principal bundle over the irrational torus with two origins $\mathbb{T}_{\alpha}^{\bullet}$.
@@
Let $\Omega_{\alpha}$, $\Omega_{\beta}$ be tiling spaces satisfying the conditions of Theorem B. Then the following conditions are equivalent:
$\Omega_{\alpha}$ and $\Omega_{\beta}$ are strong orbit equivalent
Their associated sturmian spaces $X_{\alpha},X_{\beta}$ are strong orbit equivalent
$\mathbb{T}_{\alpha}$ and $\mathbb{T}_{\beta}$ are diffeomorphic
@@
These results imply some other nice properties. On the one hand, the existence of such a bundle allows us to define and compute a fundamental group for tiling spaces thanks to the long exact sequence of homotopy groups – this appears as Corollary 1. On the other, they provide straightforward arguments for concluding some other known results about tilings belonging to the intersection of the classes of substitution and projection tilings – stated in Corollary 2.
This kind of problems have been extensively studied from different approaches as topological [26, 6, 24, 31, 35], dynamical [33, 1, 4, 3, 11], and algebraic [8, 10, 13, 14, 37, 27, 32], among many others. Our results provide both an alternative approach to previously known results as well as a basis and new techniques for further research towards an understanding of the smooth structures in tiling spaces and related objects. It is worth mentioning that there are some recent advances about the links between diffeological theory and non-commutative geometry in more general settings [22, 23, 25]. Our work could provide a series of examples for these recent advances.
The article is organized as follows. In Section 2 we introduce the basic notions of both tiling spaces and diffeology, we provide in each case several definitions and state the results that we use throughout the manuscript. In Section 3 we define and study a diffeological version of tilings spaces, we construct the first of the mentioned bundles $\hat{p}:\Omega\to\mathbb{T}_{\alpha}$ and investigate its structure groupoid, in order to prove Theorem A. Section 4 is devoted to investigate the strong orbital equivalence between tiling spaces and to the proof of the remaining results.
2 Preliminaries
@@ 2.1.1 Tilings
    Each $t_{i}$ is a translated copy of some element of a finite set of proto-tiles $\mathcal{P}=\{\tau_{1},\ldots,\tau_{k}\}$.
@@
A tiling $T$ is periodic if its symmetry group has an infinite cyclic subgroup and it is aperiodic in any other case. There are at least three methods for constructing aperiodic tilings: substitution, projection, and local rules. We will consider only the first two classes.
A substitution rule is a function from the set of proto-tiles onto the set of patches. By patch we mean a finite and connected union of tiles. It consists of a subdivision rule for the proto-tiles followed by a stretching by a linear factor $\lambda>1$, resulting in a patch of the original set of proto-tiles. By iterating a substitution rule over a finite initial configuration of tiles, arbitrarily large portions of an euclidean space can be covered and thus a tiling can be constructed.
@@ 2.1.2 Tiling spaces
By definition, $\Omega_{T}$ contains every tiling $T^{\prime}$ which is locally the same as $T$, meaning that every patch of $T^{\prime}$ may be found on $T$. In general, a tiling space $\Omega$ is a translation invariant set of tilings which is complete in the metric $d$.
@@
A substitution rule $\omega$ is called primitive if every proto-tile appears in some iteration of the substitution rule applied to any other proto-tile. Primitive substitutions generates unique $LI$ classes, thus every tiling generate the same hull and so we may speak about a tiling space $\Omega_{\omega}$ associated to $\omega$.
A tiling $T$ has finite local complexity (FLC) if for every real number $R>0$ there is a finite number of patches, up to translation, that fit inside the ball of radius $R$; and $T$ is called repetitive if for every patch $P$ of $T$ there is a radius $R>0$ such that for every $x\in\mathbb{R}^{d}$ the ball $B_{R}(x)$ contains a translated copy of $P$.
@@
Let $T$ be a $FLC$ tiling. Then $\Omega_{T}$ is compact and the following conditions are equivalent:
@@
This fairly large class of tilings produces tiling spaces with nice structures. The hulls $\Omega$ of such tilings are locally homeomorphic to cartesian products of euclidean balls $D$ and totally disconnected sets $K$ (Cantor sets under additional assumptions); they are a rich source of examples for hyperbolic dynamics, they are known to be laminated spaces [16], inverse limits of branched manifolds [2, 34, 7] and fiber bundles over tori [38, 36]. In the following paragraphs, we will state the last couple of theorems of this list and briefly explain the intuition behind them. For a deeper introduction to the topology of tiling spaces see [35].
@@
Intuition behind Theorem 2. The iteration of a substitution rule assures that arbitrarily large euclidean balls can be covered. This procedure can be encoded in the structure of an inverse limit. Regard each iteration of a substitution rule both as a finite pattern and a cylinder set, meaning that the finite pattern also represents all the tilings showing that pattern in some particular position (the canonical choice being around the origin). This cylinder set is identified with a branched manifold in the following way: consider the disjoint union of the set of prototiles and identify two of them by their borders if they appear next to each other somewhere in the tiling. Let $\Omega_{0}$ be the space obtained after the identification. It is a $CW$-complex, and it is actually a branched manifold. Then let $\Omega_{n}$ be the branched manifold obtained by applying $n$ times the substitution rule to $\Omega_{0}$ (see Fig. 1). The theorem by Anderson and Putnam [2] asserts that the tiling space $\Omega$ is topologically conjugate to the inverse limit of ${\Omega_{n},f_{n}}$ where the $f_{n}$ maps consist in forgetting the extra information from $\Omega_{n}$ to fit in $\Omega_{n-1}$. In other words, we can regard $\Omega$ as a subset of the infinite product of branched manifolds $\Pi_{n\in\mathbb{N}}\Omega_{n}$, and
$\Omega\cong\lim_{\leftarrow i}(\Omega_{n},f_{n})=\{(\bar{x}_{n})\in\Pi_{n\in\mathbb{N}}\Omega_{n}:f_{n+1}(\bar{x}_{n+1})=\bar{x}_{n},\forall n\in\mathbb{N})\}.$
@@
Figure 1: The branched 1-dimensional manifolds $\Omega_{0}$, $\Omega_{1}$ and $\Omega_{2}$ associated to the subtitution rule $a\mapsto b$, $b\mapsto ab$. Note that this image is only for illustrative purposes but it is not precisely what is needed – we are omitting here a technicality called forcing the border (see Theorem 4.3 in [2]).
@@
The idea for constructing such a bundle uses the previous result and the canonical projection $p_{0}$ from the inverse limit (which can be identified with) $\Omega$ onto the branched manifold $\Omega_{0}$, and then constructing a map from $\Omega_{0}$ which identifies each proto-tile with a torus of the corresponding dimension. The construction of the latter map relies on the fact that tiles can be deformed in order for their vertices to have rational coordinates. Such a deformation works only up to homeomorphism of tiling spaces.
Regarding the fiber, in the cases of substitution or projection tilings, it is a Cantor set, whereas for periodic tilings it is a finite collection of points.
One of the results of the current article is the construction of a variant of this bundle in the diffeological setting.
@@ 2.1.3 Equivalence of tiling spaces
  • Homeomorphism, in the usual topological sense. This is the weaker equivalence of this list. Homeomorphisms map translation orbits into translation orbits (provided the tilings have FLC), but not much more than that.
@@
Topological conjugacy is a rather rigid property among tilings. To exemplify this claim we reproduce the following example from [35]. Consider a one-dimensional tiling space $\Omega$ coming from the Fibonacci substitution $a\mapsto b$, $b\mapsto ab$, with tile lengths $|a|=1$ and $|b|=\varphi=\frac{1+\sqrt{5}}{2}$. Let $a_{n},b_{n}$ denote the patches (words) obtained by applying the substitution $n$ times to $a$ and $b$, respectively. Let $\Omega^{\prime}$ be a suspension of the Fibonacci subshift with tile lengths $|a^{\prime}|=|b^{\prime}|=(2+\varphi)/(1+\varphi)$, and let $a^{\prime}_{n}$ and $b_{n}^{\prime}$ the analogs of $a_{n},b_{n}$. The spaces $\Omega$ and $\Omega^{\prime}$ are known to be topologically conjugate by a mapping which preserves the sequences of the tiles. This happens because the differences of the lengths $|a_{n}|-|a^{\prime}_{n}|$ and $|b_{n}|-|b_{n}^{\prime}|$ go to zero as $n$ tends to infinity.
The first kind of equivalence between tiling spaces that we will encounter in sintony with diffeological equivalences is the following:
@@
Two tiling spaces $\Omega$ and $\Omega^{\prime}$ are orbit equivalent if there exist a homeomorphism $h:\Omega\to\Omega^{\prime}$ mapping the orbit $O(T)$ of $T$ in $\Omega$ onto the orbit $O(h(T))$ of $h(T)$ in $\Omega^{\prime}$ for every tiling $T\in\Omega$.
@@ 2.1.4 Geometrical and combinatorial aspects of tiling spaces
A cut-and-project set (or CPS) is a triple $(E,H,\mathcal{L})$ where $E\cong\mathbb{R}^{d}$ is the physical space of the tiling, $H$ is a locally compact abelian group called internal space, a lattice $\mathcal{L}\subset\mathbb{R}^{d}\times H$, and the canonical projections $\mathrm{pr}_{1}:\mathbb{R}^{d}\times H\to\mathbb{R}^{d}$ and $\mathrm{pr}_{2}:\mathbb{R}^{d}\times H\to H$ such that $\mathrm{pr}_{1}|_{\mathcal{L}}$ is injective and $\mathrm{pr}_{2}(\mathcal{L})$ is dense in $H$.
@@
$\Lambda(W)=\mathrm{pr}_{1}(\{y\in\mathcal{L}:\mathrm{pr}_{2}(y)\in W\}).$
@@
The tiling space of a $(\mathbb{R}^{d},\mathbb{R}^{n},\mathcal{L})$-CPS is closely related to a torus of dimension $n+d$. We explain further the case of projection from dimension 2 to 1. Consider the line $L_{\alpha}=\{y=\alpha x\}$ with slope the irrational number $\alpha$. Thus, the $(L_{\alpha},\mathbb{R},\mathbb{Z}^{2})$-CPS of canonical projection generates aperiodic tilings $T(\alpha)$ with two tiles labelled with $0,1$ of lengths $\frac{1}{\sqrt{1+\alpha^{2}}},\frac{\alpha}{\sqrt{1+\alpha^{2}}}$, respectively. Observe that each point $p$ in $L_{\alpha}$ corresponds to an element of $O(T(\alpha))$ by the relation
@@
Note that there exists an ambiguity whenever the line $L_{\alpha}$ intersects a point with integer coordinates, say $(0,0)$ for simplicity, as both projections $\pi_{2}(0,1)$ and $\pi_{2}(1,0)$ fall in the border of the canonical window $W$. Selecting either of them gives us a different tiling, and these two tilings only differ in the tiles located on their origins, so they are at maximun distance in the tiling metric.
@@
with a double solenoid $\mathcal{S}_{\alpha}^{\prime}=\pi(L_{\alpha})^{(0,1)}\cup\pi(L_{\alpha})^{(1,0)}$ and every translation orbit is represented by the projection $\pi(y=\alpha x+\rho)$ in $\mathbb{T}^{2}$, with $\rho\in(0,1)$ The double solenoid $\mathcal{S}_{\alpha}^{\prime}$ comes from the ambiguity mentioned above. See Section 1.4 in [35], and [13] for a more complete account.
@@
Given a tiling $T=\{t_{i}\}_{i\in\mathbb{Z}}$ of $\mathbb{R}$ we can identify its set of proto-tiles $\{\tau_{j}\}$ with numbers $\{a_{j}\}$ so as to obtain a sequence $S=\{s_{i}\}_{i\in\mathbb{Z}}$ which takes values in $\{a_{j}\}$. The sequence $S$ is the combinatorial aspect of the tiling $T$. This kind of sequences, endowed with the shift map
@@
The sturmian word of slope $\alpha$ and intercept $\rho$ can also be encoded as the cutting sequence111 The cutting sequence of a line $L_{\alpha}$ is constructed by writing 1 anytime the line $L_{\alpha}$ crosses an integer horizontal line between the parallels lines $x=n-1$ and $x=n$ and 0 when it doesn’t cross any integer horizontal line. of the line $y=\alpha x+\rho$. In terms of their complexity, sturmian sequences are characterized as the words containing exactly $n+1$ different subwords of lenght $n$. The literature on sturmian sequences is vast (see for example [28]). We will only focus on the types of equivalences among sturmian sequences that turn out to be relevant in our context.
@@
Two sturmian spaces $(X,\sigma)$ and $(Y,\sigma)$ are orbit equivalent if there exist a homeomorphism $h:X\to Y$ mapping the orbit $O_{\sigma}(x)$ of $x$ under $\sigma$ in $X$ onto the orbit $O_{\sigma}(h(x))$ under $\sigma$ of $h(x)$ in $Y$, for every $x\in X$.
@@
Two sturmian spaces $(X,\sigma)$ and $(Y,\sigma)$ are strong orbit equivalent if they are orbit equivalent and the cocycle map has at most one point of discontinuity.
@@
    $X_{\alpha}$ and $X_{\beta}$ are strong orbit equivalent.